# MIRACLE II: Unveiling the multi-phase gas interplay in the circumnuclear region of NGC 1365 via multi-cloud modeling

### v2

@@
We present a multi-phase analysis of the gas in the circumnuclear region ( 1.1 1.0 kpc2) of the nearby barred Seyfert 1.8 galaxy NGC 1365, observed as part of the Mid-IR Activity of Circumnuclear Line Emission (MIRACLE) program. Specifically, we combined spatially resolved spectroscopic data from JWST/MIRI, VLT/MUSE and ALMA to provide a multi-phase characterization of the ionized atomic and the warm and cold molecular gas phases. MIRI data enabled the detection of more than 40 mid-IR emission lines from ionized or warm molecular gas. Moment maps show that both cold and warm molecular gas trace the circumnuclear ring, following the rotation of the stellar disk. The ionized gas exhibits flux distributions and kinematics varying depending on the ionization potential (IP). Low-IP species ( 25 eV) mainly trace the rotating disk, while higher-IP species (up to 120 eV) trace the outflowing gas. Both [O III] and [Ne V] trace the nuclear outflow cone towards the southeast. In addition, the [Ne V] line traces the counter-cone of the outflow to the northwest, which is obscured in the optical at these circumnuclear scales and thus undetected in [O III]. Unlike optical diagnostics, spatially resolved mid-IR diagnostics reveal the key role of the active galactic nuclei (AGN) as the source of gas ionization in the central region. We derive the electron density from the [Ne V]/[Ne V] line ratio, finding a median value of (750440) cm-3, consistent with previous estimates obtained from the optical [S II] doublet. Lastly, we apply, for the first time, a fully self-consistent combination of state-of-the-art photoionization and kinematic models (HOMERUN+) to estimate the intrinsic physical outflow properties, kinematics, and energetics - overcoming the limitations of classical methods based on oversimplified assumptions. Exploiting the unprecedented synergy between JWST/MIRI and VLT/MUSE, HOMERUN allows us to simultaneously reproduce the fluxes of over 60 emission lines spanning from the optical to the mid-IR. This unique approach enables us to disentangle the physical conditions of AGN- and star formation-dominated components and robustly estimate the mass of the outflowing gas and other physical properties.
@@
NGC 1365, the subject of this work, is a great laboratory to study in detail multi-phase AGN feedback. This dusty barred spiral galaxy (SB(s)b; de Vaucouleurs et al. 1991) is located in the Fornax cluster (Jones & Jones, 1980) at a distance of 19.57 Mpc Anand et al. (2021) , corresponding to a redshift of z = 0.005457 (Bureau et al., 1996). NGC 1365 is classified as a Seyfert 1.8 galaxy (Véron-Cetty & Véron, 2006), with signature of both AGN activity and SF in its central regions.
@@
NGC 1365 has a SF rate of 16.9 M⊙ yr-1 (Lee et al., 2022), mainly within a 2 kpc radius circumnuclear ring identified across optical, radio, and IR wavelengths (Kristen et al., 1997; Forbes & Norris, 1998; Alonso-Herrero et al., 2012). The ring is associated with the inner Lindblad resonance (Lindblad et al., 1996) and shows non-circular bar-driven gas motions (Teuben et al., 1986; Sanchez, 2009), associated with SF-driven ionization (Sharp & Bland-Hawthorn, 2010; Venturi et al., 2018) and large reservoirs of molecular gas ( ; Sakamoto et al., 2007; Gao et al., 2021). Observational campaigns from the PHANGS survey (“Physics at High Angular resolution in Nearby GalaxieS”, Lee et al. 2023) conducted with the James Webb Space Telescope (JWST) and complemented by the Atacama Large Millimeter/submillimeter Array (ALMA; Wootten & Thompson 2009) have revealed that young clusters embedded in the ring drive both ionization of the surrounding medium and localized modifications to the molecular gas phase, where feedback-driven heating and dissociation decrease CO excitation and increase [C I]/CO abundance ratios in their immediate vicinity (Liu et al., 2023; Schinnerer et al., 2023).
@@
JWST is transforming our understanding of the outflow properties in the local Universe by providing an unprecedented view of the nuclear dusty regions of AGN (e.g. Hermosa Muñoz et al., 2024a; Perna et al., 2024; Zhang et al., 2024; García-Bernete et al., 2024; Davies et al., 2024; Ulivi et al., 2025; Ceci et al., 2025). This revolution is largely driven by the unique capabilities of the JWST Mid-Infrared Instrument (MIRI), including its high spatial resolution across a wide range of wavelengths (from 5 m to 28 m), enabling detailed studies of the energetics and kinematics of gas flows in obscured nuclear regions of AGN. The wide IR spectral range of JWST spanning many atomic and molecular gas transitions, combined with its spatial and spectral resolution, provides essential insights into the multi-phase circumnuclear medium. These unprecedented capabilities enable multi-phase, co-spatial analyses when combined with complementary data from ground-based facilities, such as the Multi Unit Spectroscopic Explorer (MUSE) at the ESO Very Large Telescope (VLT; Bacon et al. 2010) and ALMA, transforming our understanding of feedback mechanisms in local AGN.
@@ 2 Overview of observations and data analysis
The science observations targeted the nuclear region of NGC 1365 but were deliberately offset to include the inner part of the SE approaching outflow. As shown in the lower panels of Fig. 1, the different MIRI FoVs of each channel are not exactly centered on the nucleus, but shifted toward the SE to better capture the outflow region. This choice was motivated by the primary goal of our MIRACLE project of characterizing multi-phase gas outflows and nuclear properties. In the lower right panel of Fig. 1, we show the nuclear region of NGC 1365 in the MIRI FoV. We extracted integrated spectra from the 0.7″ radius circular apertures, and we present them in Fig. 2. The green (blue) spectrum is extracted from the circular region shown in Fig. 1 where the outflow (stellar disk) emission is dominant, as shown by the flux maps in Fig. 3. We detected more than 40 emission lines, with a signal-to-noise ratio (SNR) 5, in the spectral range covered by MIRI, mostly tracing ionized gas, with an ionization potential IP - defined as the energy required to create the relevant species - ranging from a few eV to more than 100 eV (see Table LABEL:tab:_table_list_emission_lines). Additionally, we also detected two hydrogen recombination lines, namely Pf and Hu (i.e. H I (6-5) at 7.46 m and H I (7-6) at 12.37 m, respectively) and seven H2 pure-rotational transitions, from 0-0 S(7) to 0-0 S(1).
As shown in Fig. 2, the MIRI spectra also exhibit at least 15 bright PAH features in the wavelength range between 5 - 18 m, which we will investigate in a dedicated forthcoming paper. To obtain these spectra, for the first time we combined the emission from all 12 MIRI datacubes, taking into account the different pixel sizes and ensuring a proper flux conservation among different bands by applying a scaling factor to align the flux levels between adjacent bands. A detailed description of this novel procedure is presented in the Appendix B.
In Appendix C.1, we analyzed the MIRI MRS datacubes using a customized python script to enhance the SNR with spatial smoothing, subtract the local continuum around all the brightest emission lines and then fit them spaxel by spaxel. The fitted emission lines are listed in Table LABEL:tab:_table_list_emission_lines.
@@ 3.1 Multi-phase gas kinematics
For the kinematic analysis, we focused on emission lines within MIRI Ch3 (0.2″/pixel in a 6.6″ 7.7″ FoV, corresponding to 23 pc/pixel over a 770 900 pc2 FoV) because of the good balance between spatial resolution and field coverage. Moreover, the Ch3 emission lines are among the brightest detected in our data, allowing us to present spatially resolved results representative of the ionized and molecular emission, with sufficient SNR across the entire FoV. Although in this analysis we focus on the FoV of Ch3, which also includes a fraction of disk gas illuminated by the AGN and not part of the outflow, it would be more accurate to refer to this region as the ionization cone rather than calling it outflow. However, we will assume in the following that all the gas in the ionization cone is in outflow.
Figure 3 shows the moment maps of [Ne V], [Ne II], H2 0-0 S(1), and CO(3-2), obtained from the emission line fitting routines presented in Appendices C.1 and C.3. The [Ne V] seems to follow the direction of the ionized outflow, that is, the SE-NW direction (Venturi et al. 2018), as expected from species with high IP (see Table LABEL:tab:_table_list_emission_lines). Its velocity field is also consistent with the [O III] kinematics, which is blueshifted in the SE and redshifted in the NW. In the SE (NW) cone, we estimate gas projected velocities up to -120 km s-1 (+100 km s-1). The velocity dispersion map of [Ne V] shows a donut-shaped structure in SE direction of the nucleus (highlighted with contours in Fig. 4), at a projected distance of 2.5″ ( 300 pc), with a peak of 170 km s-1.
@@
The CO(3-2) moment maps shown in Fig. 3 are in agreement with the H2 0-0 S(1) kinematics, that is tracing the galaxy rotating disk. A similar CO velocity field was reported by Liu et al. (2023). This finding, which is consistent with the lack of any molecular outflow in the circumnuclear region (see also Combes et al., 2019), is in contrast to the CO (1-0) wind reported by Gao et al. (2021). Their outflow interpretation was based on velocity residuals from a rotating disk model on larger spatial scales, which resemble those observed in H and are more likely due to non-circular inflowing motions along the bar (see their Fig. 6). Notably, the cold gas traced by the CO(3-2) does not show the -enhanced region south of the nucleus as the H2, although both emission lines map the molecular gas. This suggests that the elevated velocity dispersion in H2 may trace non-circular motions, possibly related to local turbulence or streaming motions induced by the AGN, rather than a genuine outflow. In such an environment, the cold molecular gas traced by CO may not survive due to stronger dissociation or heating, leaving only the warm phase observable in this disturbed region.
The moment maps of the other emission lines listed in Table LABEL:tab:_table_list_emission_lines are shown in Appendix H, divided into molecular gas, recombination lines, and species with low ( 25 eV), medium ( 25 and 80 eV), and high ( 80 eV) IP. Grouping such emission lines by IP or gas phase leads to interesting typical patterns. All the H2 lines share kinematic features similar to those of H2 0-0 S(1) in Fig. 3, with a rotational disk kinematics and a high- blob to the South of the nucleus. The low IP lines show a velocity gradient consistent with a rotating disk, with an amplitude of 110 km s-1, like [Ne II]. The presence of the high- region observed in H2 is not present in all lines. From the flux maps, we observed that these species are located mainly in the ‘Mid-East’ region (Liu et al., 2023). The moment maps of the high-IP species are similar to those of [Ne V] (see Fig. 3). In particular, the donut-shaped structure is well defined in the velocity dispersion maps, although in the Ch4 the spatial resolution is lower and the shape is not well recognizable. The medium-IP lines have intermediate properties compared to the previous two groups: the velocity gradient is similar to that of the low-IP lines, but the velocity dispersion maps show a high-velocity dispersion region in correspondence of the donut shape.
@@ 3.3 Resolved diagnostic diagrams for gas excitation
Diagnostic diagrams featuring mid-IR emission line ratios are widely used to investigate the nature of the gas ionizing sources (e.g. Hao et al., 2009; Weaver et al., 2010; Inami et al., 2013; Richardson et al., 2022; Martínez-Paredes et al., 2023; Feltre et al., 2023; Garofali et al., 2024; Mingozzi et al., 2025). In particular, to map the gas excitation within the circumnuclear region of NGC 1365 we exploited the mid-IR line ratios recently explored by Feltre et al. (2023) within local AGN, exploiting Spitzer integrated spectra including NGC 1365. To account for the different FoVs, pixel sizes, and point spread functions (PSFs) of each sub-channel, we corrected each flux map of the mentioned emission lines accordingly before computing the ratios. Specifically, we adopted the smallest FoV (6.6″ 7.7″, i.e. Ch3), the largest pixel size (0.35″ from Ch4), and convolved all images to the worst PSF (i.e. 1.0″, the one at the reddest wavelength of m). Additionally, as detailed in Appendix B, we applied multiplicative correction factors to account for flux discontinuities between the datacubes (see Fig. 11).
@@
In the upper left panel of Fig. 5, we show the contours of line ratio model predictions for AGN (Feltre et al., 2023), SF galaxies (Gutkin et al., 2016), and shocks (Alarie & Morisset, 2019) for the [O IV]m/[Ne II]m and [Ne III]m/[Ne II]m line ratios. We include all available model grids, applying only a metallicity cut of (see Feltre et al. 2023 for further details on the models). We overlay our data on model predictions, where each pixel is color-coded based on its position in the corresponding diagnostic diagram. This diagram shows that the main ionization source in the circumnuclear region of NGC 1365 can either be shocks or AGN, while no evidence for SF excitation is detected. According to Fig. 5 of Feltre et al. (2023), adding fractional contributions from SF and shock models to an AGN model will shift the points towards the bottom-left in the diagram. Therefore, we classify as shock- or AGN-excited only those spaxels that fall entirely within a single predicted region. Data points located in overlapping or unclassified regions cannot be uniquely attributed to a single excitation mechanism. To better visualize the spatial distribution and morphology of these regions, we produce diagnostic maps using the same color-coding of the diagrams, as shown in the lower panels of Fig. 5.
@@
In the central column of Fig. 5, we show the [Ne III]m/[Ne II]m versus [Ne V]m/[Ne II]m diagnostic diagram. Similarly to the [O IV] diagram, the data points are separated into shock- and/or AGN-excited models. The [O IV] and [Ne V] excitation maps reveal similar spatial structures: an extended region dominated by AGN and combined shock/AGN excitation, aligned with the disk major axis and delimited on either side by shock-dominated zones. Notably, the eastern shock front coincides with the edge of the circumnuclear ring traced by CO and H (see Fig. 3 and Venturi et al., 2018), which may indicate that the outflow, emerging from the disk, shocks the material of the circumnuclear ring along its path. In contrast, the centrally elongated feature may reflect direct AGN ionization within the disk.
@@
Exploiting optical MUSE data, we compare the mid-IR diagnostic diagrams with the optical [S II] BPT diagram (Kewley et al., 2006, 2013) shown in the right column of Fig. 5. Moreover, we also considered the [N II] and [O I] BPT diagrams, finding similar results (see Fig. 5 in Venturi et al., 2018 and Appendix D3 in Mingozzi et al., 2019 for these diagrams at large scale). As found by Venturi et al. (2018), we notice that the circumnuclear SF ring seen in H and in CO (see Fig. 1) is not fully dominated by SF in the BPT diagrams. Instead, as visible from the [S II] spatial distribution, certain regions in the SE of the nucleus show signatures of AGN excitation. Venturi et al. (2018) conclude that this is likely due to the superposition along the line of sight of the AGN-ionized cone and the SF ring. The main difference between the optical BPT and the mid-IR diagnostic diagrams we presented lies in their ability to trace the true ionization source in dusty environments. In the nuclear region, the optical BPT appears to be dominated by star formation due to strong dust obscuration, as indicated by the Balmer decrement (see Appendix D and Fig. 12). In contrast, mid-IR diagnostics can penetrate the dust screen and clearly reveal AGN ionization at the nucleus. This explains why, in the optical, AGN ionization cones are only visible on larger spatial scales, as previously shown in Venturi et al. (2018) and Mingozzi et al. (2019). Another notable difference is that the region classified as shock-dominated in the mid-IR diagrams appears to be excited by the AGN radiation according to the optical diagnostic diagram. This discrepancy likely reflects the different depths probed by optical and mid-IR emission, with mid-IR lines tracing higher ionization regions deeper into the disk, where shocks from the interaction between the outflow and the disk are expected to occur.

### v3

@@
We present a multi-phase analysis of the gas in the circumnuclear region ( 0.90.9 kpc2) of the nearby barred Seyfert 1.8 galaxy NGC 1365, observed as part of the Mid-IR Activity of Circumnuclear Line Emission (MIRACLE) program. Specifically, we combined spatially resolved spectroscopic data from JWST/MIRI, VLT/MUSE and ALMA to provide a multi-phase characterization of the ionized atomic and the warm and cold molecular gas phases. MIRI data enabled the detection of more than 40 mid-IR emission lines from ionized or warm molecular gas. Moment maps show that both cold and warm molecular gas trace the circumnuclear ring, following the rotation of the stellar disk. The ionized gas exhibits flux distributions and kinematics varying depending on the ionization potential (IP). Low-IP species ( 25 eV) mainly trace the rotating disk, while higher-IP species (up to 120 eV) trace the outflowing gas. Both [O III] and [Ne V] trace the nuclear outflow cone towards the southeast. In addition, the [Ne V] line traces the counter-cone of the outflow to the northwest, which is obscured in the optical at these circumnuclear scales and thus undetected in [O III]. Unlike optical diagnostics, spatially resolved mid-IR diagnostics reveal the key role of the active galactic nuclei (AGN) as the source of gas ionization in the central region. We derive the electron density from the [Ne V]/[Ne V] line ratio, finding a median value of (750440) cm-3, consistent with previous estimates obtained from the optical [S II] doublet. Lastly, we apply, for the first time, a fully self-consistent combination of state-of-the-art photoionization and kinematic models (HOMERUN+) to estimate the intrinsic physical outflow properties, kinematics, and energetics - overcoming the limitations of classical methods based on oversimplified assumptions. Exploiting the unprecedented synergy between JWST/MIRI and VLT/MUSE, HOMERUN allows us to simultaneously reproduce the fluxes of over 60 emission lines spanning from the optical to the mid-IR. This unique approach enables us to disentangle the physical conditions of AGN- and star formation-dominated components and robustly estimate the mass of the outflowing gas and other physical properties.
@@
NGC 1365, the subject of this work, is a great laboratory to study in detail multi-phase AGN feedback. This dusty barred spiral galaxy (SB(s)b; de Vaucouleurs et al. 1991) is located in the Fornax cluster (Jones & Jones, 1980) at a distance of 19.57 Mpc (Jacobs et al., 2009; Anand et al., 2021) (1″ 95 pc). With a redshift of z = 0.005457 (Bureau et al., 1996), NGC 1365 is classified as a Seyfert 1.8 galaxy (Véron-Cetty & Véron, 2006), with signature of both AGN activity and SF in its central regions.
@@
NGC 1365 has a SF rate of 16.9 M⊙ yr-1 (Lee et al., 2022), mainly within a 2 k radius circumnuclear ring identified across optical, radio, and IR wavelengths (Kristen et al., 1997; Forbes & Norris, 1998; Alonso-Herrero et al., 2012). The ring is associated with the inner Lindblad resonance (Lindblad et al., 1996) and shows non-circular bar-driven gas motions (Teuben et al., 1986; Sanchez, 2009), associated with SF-driven ionization (Sharp & Bland-Hawthorn, 2010; Venturi et al., 2018) and large reservoirs of molecular gas ( ; Sakamoto et al., 2007; Gao et al., 2021). Observational campaigns from the PHANGS survey (“Physics at High Angular resolution in Nearby GalaxieS”, Lee et al. 2023) conducted with the James Webb Space Telescope (JWST) and complemented by the Atacama Large Millimeter/submillimeter Array (ALMA; Wootten & Thompson 2009) have revealed that young clusters embedded in the ring drive both ionization of the surrounding medium and localized modifications to the molecular gas phase, where feedback-driven heating and dissociation decrease CO excitation and increase [C I]/CO abundance ratios in their immediate vicinity (Liu et al., 2023; Schinnerer et al., 2023).
@@
JWST is transforming our understanding of the outflow properties in the local Universe by providing an unprecedented view of the nuclear dusty regions of AGN (e.g. Hermosa Muñoz et al., 2024a; Perna et al., 2024; Zhang et al., 2024; García-Bernete et al., 2024; Davies et al., 2024; Ulivi et al., 2025; Ceci et al., 2025). This revolution is largely driven by the unique capabilities of the JWST Mid-Infrared Instrument (MIRI; Wright et al., 2015, 2023), including its high spatial resolution across a wide range of wavelengths (from 5 m to 28 m), enabling detailed studies of the energetics and kinematics of gas flows in obscured nuclear regions of AGN. The wide IR spectral range of JWST spanning many atomic and molecular gas transitions, combined with its spatial and spectral resolution, provides essential insights into the multi-phase circumnuclear medium. These unprecedented capabilities enable multi-phase, co-spatial analyses when combined with complementary data from ground-based facilities, such as the Multi Unit Spectroscopic Explorer (MUSE) at the ESO Very Large Telescope (VLT; Bacon et al. 2010) and ALMA, transforming our understanding of feedback mechanisms in local AGN.
@@ 2 Overview of observations and data analysis
The science observations targeted the nuclear region of NGC 1365 but were deliberately offset to include the inner part of the SE approaching outflow. As shown in the lower panels of Fig. 1, the different MIRI FoVs of each channel are not exactly centered on the nucleus, but shifted toward the SE to better capture the outflow region. This choice was motivated by the primary goal of our MIRACLE project of characterizing multi-phase gas outflows and nuclear properties. In the lower right panel of Fig. 1, we show the nuclear region of NGC 1365 in the MIRI FoV. We extracted integrated spectra from the 0.7″ radius circular apertures, and we present them in Fig. 2. The green (blue) spectrum is extracted from the circular region shown in Fig. 1 where the outflow (stellar disk) emission is dominant, as is shown by the flux maps in Fig. 3. Note that the stellar disk spectrum is contaminated by outflow emission, as indicated by the presence of high-excitation lines. We detected more than 40 emission lines, with a signal-to-noise ratio (S/N) 5, in the spectral range covered by MIRI, mostly tracing ionized gas, with an ionization potential IP - defined as the energy required to create the relevant species - ranging from a few eV to more than 100 eV (see Table LABEL:tab:_table_list_emission_lines). Additionally, we also detected two hydrogen recombination lines, namely Pf and Hu (i.e. H I (6-5) at 7.46 m and H I (7-6) at 12.37 m, respectively) and seven H2 pure-rotational transitions, from 0-0 S(7) to 0-0 S(1).
As shown in Fig. 2, the MIRI spectra also exhibit at least 15 bright PAH features in the wavelength range between 5 - 18 m, which we will investigate in a dedicated forthcoming paper. To obtain these spectra, for the first time we combined the emission from all 12 MIRI datacubes, taking into account the different pixel sizes and ensuring a proper flux conservation among different bands by applying a scaling factor to align the flux levels between adjacent bands. A detailed description of this novel procedure is presented in the Appendix B. This algorithm, written in Python, is available for download111https://github.com/matteo-ceci/JWST-MIRI_stitching_code.
In Appendix C.1, we analyzed the MIRI MRS datacubes using a customized python script to enhance the S/N with spatial smoothing, subtract the local continuum around all the brightest emission lines and then fit them spaxel by spaxel. The fitted emission lines are listed in Table LABEL:tab:_table_list_emission_lines.
@@ 3.1 Multi-phase gas kinematics
For the kinematic analysis, we focused on emission lines within MIRI Ch3 (0.2″/pixel in a 6.6″ 7.7″ FoV, corresponding to 19 pc/pixel over a 620 730 pc2 FoV) because of the good balance between spatial resolution and field coverage. Moreover, the Ch3 emission lines are among the brightest detected in our data, allowing us to present spatially resolved results representative of the ionized and molecular emission, with sufficient S/N across the entire FoV. Although in this analysis we focus on the FoV of Ch3, which also includes a fraction of disk gas illuminated by the AGN and not part of the outflow, it would be more accurate to refer to this region as the ionization cone rather than calling it outflow. However, we will assume in the following that all the gas in the ionization cone is in outflow.
Figure 3 shows the moment maps of [Ne V], [Ne II], H2 S(1), and CO(3-2), obtained from the emission line fitting routines presented in Appendices C.1 and C.3. The [Ne V] seems to follow the direction of the ionized outflow, that is, the SE-NW direction (Venturi et al. 2018), as expected from species with high-IP (see Table LABEL:tab:_table_list_emission_lines). Its velocity field is also consistent with the [O III] kinematics, which is blueshifted in the SE and redshifted in the NW. In the SE (NW) cone, we estimate gas projected velocities up to -120 km s-1 (+100 km s-1). The velocity dispersion map of [Ne V] shows a donut-shaped structure in SE direction of the nucleus (highlighted with contours in Fig. 4), at a projected distance of 2.5″ ( 240 pc), with a peak of 170 km s-1.
@@
The CO(3-2) moment maps shown in Fig. 3 are in agreement with the H2 S(1) kinematics, that is tracing the galaxy rotating disk. A similar CO velocity field was reported by Liu et al. (2023). This finding, which is consistent with the lack of any molecular outflow in the circumnuclear region (see also Combes et al., 2019), is in contrast to the CO (1-0) wind reported by Gao et al. (2021). Their outflow interpretation was based on velocity residuals from a rotating disk model on larger spatial scales, which resemble those observed in H and are more likely due to non-circular inflowing motions along the bar (see their Fig. 6). Notably, the cold gas traced by the CO(3-2) does not show the -enhanced region south of the nucleus as the H2, although both emission lines map the molecular gas. This suggests that the elevated velocity dispersion in H2 may trace non-circular motions, possibly related to local turbulence or streaming motions induced by the AGN, rather than a genuine outflow. In such an environment, the cold molecular gas traced by CO may not survive due to stronger dissociation or heating, leaving only the warm phase observable in this disturbed region.
The moment maps of the other emission lines listed in Table LABEL:tab:_table_list_emission_lines are shown in the online material, divided into molecular gas, recombination lines, and species with low ( 25 eV), medium ( 25 and 54 eV), and high ( 54 eV) IP. The thresholds are based on the ionization potentials of He I (25 eV) and He II (54 eV). Note that while some lines appear to lack emission in the nuclear region of the galaxy in the moment maps, this is due to low S/N where the spectra are dominated by strong continuum emission, rather than a genuine absence of emission. Grouping such emission lines by IP or gas phase leads to interesting typical patterns. All the H2 lines share kinematic features similar to those of H2 S(1) in Fig. 3, with a rotational disk kinematics and a high- blob to the South of the nucleus. The low-IP lines, as well as the recombination lines, show a velocity gradient consistent with a rotating disk, with an amplitude of 110 km s-1, like [Ne II]. The high- region observed in H2 is not present in all the low-IP and recombination lines. From the flux maps, we observed that these species are located mainly in the ‘Mid-East’ region (Liu et al., 2023). The moment maps of the high-IP species are similar to those of [Ne V] (see Fig. 3). In particular, the donut-shaped structure is well defined in the velocity dispersion maps, although in the Ch4 the spatial resolution is lower and the shape is not well recognizable. The medium-IP lines have intermediate properties compared to the previous two groups: the velocity gradient is similar to that of the low-IP lines, but the velocity dispersion maps show a high-velocity dispersion region in correspondence of the donut shape.
@@ 3.3 Resolved diagnostic diagrams for gas excitation
Diagnostic diagrams featuring mid-IR emission line ratios are widely used to investigate the nature of the gas ionizing sources (e.g. Hao et al., 2009; Weaver et al., 2010; Inami et al., 2013; Richardson et al., 2022; Martínez-Paredes et al., 2023; Feltre et al., 2023; Garofali et al., 2024; Mingozzi et al., 2025). In particular, to investigate the gas excitation in the circumnuclear region of NGC 1365, we employed mid-IR line ratios recently proposed by Feltre et al. (2023), who analyzed integrated nuclear spectra of 42 local Seyfert galaxies observed with Spitzer, including NGC 1365. To ensure consistent comparisons between line ratios, we corrected the flux maps of each emission line for differences in FoV, pixel scale, and point spread functions (PSF). In practice, for each diagnostic diagram we adopted the FoV, pixel size, and PSF corresponding to the worst case among the lines involved, as detailed below. Additionally, as detailed in Appendix B, we applied multiplicative correction factors to account for flux discontinuities between the datacubes (see Fig. 11).
@@
In the upper left panel of Fig. 5, we show the contours of line ratio model predictions for AGN (Feltre et al., 2023), SF galaxies (Gutkin et al., 2016), and shocks (Alarie & Morisset, 2019) for the [O IV]m/[Ne II]m and [Ne III]m/[Ne II]m line ratios. For this diagnostic diagram, we adopted the smallest FoV (6.6″ 7.7″, i.e. Ch3), the largest pixel size (0.35″ from Ch4), and convolved all images to the worst PSF (i.e. 1.0″, the one at the reddest wavelength of m). We include all available model grids, applying only a metallicity cut of (see Feltre et al. 2023 for further details on the models). We overlay our data on model predictions, where each pixel is color-coded based on its position in the corresponding diagnostic diagram. This diagram shows that the main ionization source in the circumnuclear region of NGC 1365 can either be shocks or AGN, while no evidence for SF excitation is detected. According to Fig. 5 of Feltre et al. (2023), adding fractional contributions from SF and shock models to an AGN model will shift the points towards the bottom-left in the diagram. Therefore, we classify as shock- or AGN-excited only those spaxels that fall entirely within a single predicted region. Data points located in overlapping or unclassified regions cannot be uniquely attributed to a single excitation mechanism. To better visualize the spatial distribution and morphology of these regions, we produce diagnostic maps using the same color-coding of the diagrams, as shown in the lower panels of Fig. 5.
@@
In the central column of Fig. 5, we show the [Ne III]m/[Ne II]m versus [Ne V]m/[Ne II]m diagnostic diagram. In this case, all lines fall within the same channel. Therefore, only the PSF convolution was required, while the FoV and pixel scale were already consistent across the maps. Similarly to the [O IV] diagram, the data points are separated into shock- and/or AGN-excited models. The [O IV] and [Ne V] excitation maps reveal similar spatial structures: an extended region dominated by AGN and combined shock/AGN excitation, aligned with the disk major axis and delimited on either side by shock-dominated zones. Notably, the eastern shock front coincides with the edge of the circumnuclear ring traced by CO and H (see Fig. 3 and Venturi et al., 2018), which may indicate that the outflow, emerging from the disk, interacts with the material of the circumnuclear ring, producing shocks along its path. In contrast, the centrally elongated feature may reflect direct AGN ionization within the disk.
However, Laor (1998) showed that photoionizing shocks are extremely inefficient in powering the narrow line emission in AGN. Furthermore, as noted by Feltre et al. (2023), the predictions from shock models largely overlap with both AGN and star-formation grids in diagnostic diagrams. For these reasons, in Section 3.5 we consider only AGN and SF grids to model the optical and mid-IR emission lines.
@@
Exploiting optical MUSE data, we compare the mid-IR diagnostic diagrams with the optical [S II] BPT diagram (Kewley et al., 2006, 2013) shown in the right column of Fig. 5. Moreover, we also considered the [N II] and [O I] BPT diagrams, finding similar results (see Fig. 5 in Venturi et al., 2018 and Appendix D3 in Mingozzi et al., 2019 for these diagrams at large scale). As found by Venturi et al. (2018), we notice that the circumnuclear SF ring seen in H and in CO (see Fig. 1) is not fully dominated by SF in the BPT diagrams. Instead, as visible from the [S II] spatial distribution, certain regions in the SE of the nucleus show signatures of AGN excitation. Venturi et al. (2018) conclude that this is likely due to the superposition along the line of sight of the AGN-ionized cone and the SF ring. The main difference between the optical BPT and the mid-IR diagnostic diagrams lies in their sensitivity to trace the true ionization source in dusty environments. In the nuclear region, the optical BPT appears to be dominated by star formation due to strong dust obscuration, as indicated by the Balmer decrement (see Appendix D and Fig. 12). In contrast, mid-IR diagnostics can penetrate the dust screen and clearly reveal AGN ionization at the nucleus. This explains why, in the optical, AGN ionization cones are only visible on larger spatial scales, as previously shown in Venturi et al. (2018) and Mingozzi et al. (2019). Another notable difference is that the region classified as shock-dominated in the mid-IR diagrams appears to be excited by the AGN radiation according to the optical diagnostic diagram. This discrepancy likely reflects the different depths probed by optical and mid-IR emission, with mid-IR lines tracing higher ionization regions deeper into the disk, where shocks from the interaction between the outflow and the disk are expected to occur.